a strong current of trivalent gaseous boron passed thru a germanium crystal decreases d density of crystal due to part replacement of germanium by boron and due to interstitial vacancies created by missing germanium atoms.In one such xperiment one gram of germanium is taken and the boron atoms r found to be 150ppm by wt,when density of germanium crystal decreases by 4%.calculate d % of missing vacancies due to germanium which r filled up by boron atoms???


 


pls explain each step as well.


thank u.

Grade:12

1 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi Simmi singh,

1 g Germanium------- 150 ppm= 1.5 * 10^-4 g Boron

4% decrease---- now weight is 0.96 g(Ge + B)----0.04 g missing

0.04 missing-------- (0.04-0.00015)g Ge missing------(0.04-0.00015)/72.61 moles Ge missing=5.488*10^-4 moles

Boron replaced= 1.5*10^-4 g = (1.5*10^-4)/10 moles

                equvalently replaced = 3/4((1.5*10^-4)/10) moles of Ge    ( Valency of B is 3 and Ge is 4)

                                              = 0.1125*10^-4 moles of Ge

% of Missing vacancies filled by B atoms = 0.1125*100/5.488 = 2.05 %